OpenStudy (anonymous):

"If it is an identity or has no soultion, please write whitch it is." -(-5x=4)=3x-10

OpenStudy (gorv):

take minus inside 5x-4=3x-10 subtract 3x from both side \ 5x-3x-4=-10 2x-4=-10 add 4 on both side

OpenStudy (gorv):

and divide by 2

OpenStudy (anonymous):

Thank you ^_^

OpenStudy (gorv):

welcome to os btw :)